OpenStudy (anonymous):

find partial derivative Hx and Hy of H(x,y)=(y^2+1)e^x

OpenStudy (cruffo):

can you explain how you find the partial derivative with respect to x? Short sentance...

OpenStudy (anonymous):

you basically do derivative just with respect to x and treating y as a constant

zepdrix (zepdrix):

So Mproof, Hmm If you take partials, you'll be treating the OTHER variable as a constant, meaning you won't have the product rule as it might seem at first glance. Does that help? :O

OpenStudy (anonymous):

like 3x^2y+2 fx=6xy and Fy is 3x^2

OpenStudy (cruffo):

right. So what is confusing you about this problem?

OpenStudy (anonymous):

would Hx be 0?

OpenStudy (cruffo):

no. What is the regular derivative of f(x) = e^x?

OpenStudy (anonymous):

same

zepdrix (zepdrix):

Think of the equation as Ce^x when taking the partial WRT x. Maybe that will help :)

OpenStudy (anonymous):

but don't you have to take a derivative of the (y^2+1) with respect to x?

zepdrix (zepdrix):

no, thats just a constant attached to e^x :d

zepdrix (zepdrix):

Maybe one thing you can do to convince yourself is, distribute the e^x to each term in the brackets. Then think about what you have :o

OpenStudy (anonymous):

so the derivative with respect to x will be the same as the given problem

zepdrix (zepdrix):

ya :) good

OpenStudy (anonymous):

ooo I get it

OpenStudy (anonymous):

with respect to Y would it be 2ye^x?

zepdrix (zepdrix):

(y^2 + 1)e^x = y^2 e^x + e^x Hy = 2y e^x + 0 Yes, very good ^^